OpenStudy (anonymous):

I really need help plz What is the slope of the line that passes through the points (–5, 3) and (7, 5)? A.–1 B. 1/6 C.1 D.6

hartnn (hartnn):

The slope of the line through points (x1,y1) and (x2,y2) is given by : \(\huge m=\frac{y_1-y_2}{x_1-x_2}\) now,just put the values and find m.

OpenStudy (anonymous):

hi

hartnn (hartnn):

@Sora78 \(\Huge \mathcal{\text{Welcome To OpenStudy}\ddot\smile} \)

OpenStudy (anonymous):

@hartnn thanks for the help :) i appreciate it

hartnn (hartnn):

you're most welcome ^_^

OpenStudy (anonymous):

@hartnn if my math is correct and it usually isn't 6 would be correct is that right

hartnn (hartnn):

numerator = y1-y2 = 5-3 = 2 u sure its 6 and not 1/6 ?

OpenStudy (anonymous):

as i said usually not right i re figured it 1/6

hartnn (hartnn):

good! :) and yes its 1/6 :)

OpenStudy (anonymous):

you can use the formula of slope..